OpenStudy (anonymous):

Find the amplitude , period phase shift and vertical shift for each function. provide a table of values and then sketch the graph y=3sin(x-pi/2)

OpenStudy (raffle_snaffle):

amplitude is determined by the # in front ot the sin

OpenStudy (ybarrap):

Here's the anatomy of a periodic function: $$ y=A\sin(\cfrac{2\pi}{T}t+\theta) + b $$ Where \(A\) is the amplitude, \(T\) is the period, \(b\) is the vertical shift and \(\theta\) is the phase.